Using four 4s, no other numbers, and as many mathematical operations as you like, create mathematical expressions that equal the integers from 1 on. How high can you go?
Answer by vleith(2517)

1 = (44/44)
2 = (4/4) + (4/4)
3 = (4+4+4)/4
4 = sqrt(4*4) *(4/4)
5 = sqrt(4*4) + (4/4)
6 = 4 + (4+4)/4
7 = (44/4) - 4
8 = (4*4)/(4+4)
9 = (4+4) + (4/4)